OpenStudy (anonymous):

1. Use the slope formula to find the slope of the line that passes through each pair of points. a. (-5, 7) and (0, 0) b. (-3, -5) and (3, 11) c. (3, -5) and (-2, 9) d. (-5, 7) and (-5, 11) e. (9, 9) and (-9, -9) f. (3, 5) and (-2, 7) g. (2.5, 3) and (8, 3.5)

OpenStudy (reemii):

you wil appply the formula \(\frac{\Delta y}{\Delta x}\) for each pair of points. a. \(\frac{7-0}{-5-0}=-\frac75\). etc. we can check the answers.

OpenStudy (anonymous):

idk what that is ?

hero (hero):

\[m = \dfrac{\Delta y}{\Delta x} = \dfrac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

not really